# Position: Simulating Society Requires Simulating Thought

### v2

@@ Abstract
Simulating society with large language models (LLMs), we argue, requires more than generating plausible behavior; it demands cognitively grounded reasoning that is structured, revisable, and traceable. LLM-based agents are increasingly used to emulate individual and group behavior, primarily through prompting and supervised fine-tuning. Yet they often lack internal coherence, causal reasoning, and belief traceability, making them unreliable for simulating how people reason, deliberate, and respond to interventions.
@@ 1 Introduction
Over the past two years, LLMs have increasingly become dominant tools for simulating human behavior across language [1, 2], vision [3] and decision-making domains [4].In the field of social simulation, LLMs are now commonly used to emulate public opinions, stakeholder interactions, and policy responses under diverse scenarios [5, 6].
Despite these advances, existing models remain far from cognitively intelligent. They often exhibit shallow reasoning, frequent hallucinations, and limited understanding of commonsense causal relationships in socially-relevant topics such as upzoning, surveillance, or healthcare access [7, 8].
In addition, these models are difficult to adapt to new domains. Fine-tuning LLMs for context-specific simulation requires substantial computational resources and access to large, high-quality datasets, both of which are often scarce in real-world policy settings. Yet effective social simulation demands precisely the opposite: the ability to represent dynamically evolving stakeholder reasoning, grounded in timely and localized information [9, 10].
Moreover, current LLM-based agents remain largely opaque in their internal reasoning under contextual change. It is difficult to determine how much a given policy change or contextual shift influences the agent’s opinions or decisions—let alone why. Most existing systems operate as black boxes, lacking interpretable reasoning traces that reveal how conclusions are formed. This is a critical limitation in the social simulation domain, where understanding how individuals arrive at particular judgments is often more important than the judgments themselves.
While post-hoc analysis of outputs is common, it is insufficient for modeling complex, multi-stakeholder reasoning. Without transparent internal structures, agents cannot support diagnostic explanation, causal attribution, or meaningful intervention—all of which are essential for interpreting public responses to policy and for building trustworthy simulations.
@@
These limitations call for a shift in how we conceptualize generative social simulation—not as behavior mimicry, but as cognitive modeling. This paper takes up that call. Specifically, we propose leveraging ideas from Theory of Mind (ToM) and cognitive science to extract and simulate structured reasoning templates—what we term reasoning traces—rather than simply mimicking human tone or persona [11, 12].
@@
Such modularity is the cornerstone of human cognition, where reasoning is built from reusable, revisable fragments across contexts [14, 15]. It also makes simulation more faithful: interactions between agents, or between agents and environment dynamics, can be represented as separable, inspectable modules—enabling structured simulation at both micro and macro levels [16]. Moreover, by reusing modular reasoning components rather than re-generating full-context reasoning each time, this approach significantly reduces the token budget required during simulation and inference—making it both more interpretable and more computationally efficient.
In this paper, we advocate moving beyond output-level alignment toward aligning the internal reasoning traces of generative agents. Capturing the causal, compositional, and revisable structure of belief formation—what we call reasoning fidelity—is essential to building cognitively faithful agents that simulate not only what people say, but how they think.
@@
  • Present a methodology for extracting and simulating belief structures from natural language, enabling interpretability, counterfactual reasoning, and domain transfer;
  • Illustrate how current approaches fall short by producing outputs that appear coherent but lack internal consistency, adaptability, or traceability.
@@ 2 Social Simulation: Opportunities and Gaps
Social simulation has emerged as a high-impact use case for LLMs, particularly in policy modeling, urban planning, and behavior forecasting. Traditionally, social science relies on surveys, experiments, and fieldwork to understand individual and group behavior. While effective, these methods are expensive, hard to scale, and often face ethical and logistical challenges. The rise of LLM-driven agents offers a promising alternative—capable of simulating human responses across a wide range of scenarios, roles, and interventions.
@@
Beyond single-agent settings, multi-agent simulations further extend this potential by modeling the interactions, conflicts, and consensus dynamics among synthetic populations. However, faithfully simulating social processes introduces several critical requirements:
@@
    Fidelity: Simulated agents must remain sensitive to environmental and policy contexts, updating their beliefs and responses accordingly [25, 26].
@@
    Diversity: Outputs should reflect the heterogeneity of real populations—avoiding “flattening” effects where different demographics collapse into a single voice [25, 27].
@@
    Scalability: Simulations must scale to the population level to support distributional analysis of public opinion and policy impact [28, 29].
  4. 4.
    Generalization: Models should extrapolate to unseen cases by reusing previously learned reasoning motifs across domains [26, 27].
  5. 5.
    Consistency: A simulated agent should maintain a coherent reasoning trace across different scenarios, with beliefs that evolve predictably rather than being reset at each prompt [29].
While many recent works aim to create more “human-like” agents, few clearly define what this term means—or how such fidelity is to be evaluated. Most existing approaches focus on behavior-level plausibility, often relying on post-hoc justification rather than simulating the underlying structure of belief formation and revision.
In this work, we take a complementary approach by grounding agent simulation in cognitive theory. We argue that modeling reasoning traces—explicit, modular structures of how beliefs are formed and revised—enables more faithful, transparent, and generalizable social simulations. Our aim is not only to guide simulator design but also to provide clearer standards for evaluating claims of human-likeness in generative agents. While our primary focus is social simulation, we believe our approach contributes to a broader shift toward cognitively grounded agent modeling in human-AI systems.
@@
3 Problem Statement: Behavioral Plausibility is Not Enough
@@
Most existing efforts to align LLM-based agents focus on outputs [30]: Do they express coherent stances, mimic preferences, or engage in natural-sounding conversations? This behavior-centric view is reinforced by popular techniques such as reinforcement learning from human feedback (RLHF) [31, 32], persona prompting [33], and chain-of-thought (CoT) generation [17, 34]. These methods optimize for plausibility—not for structural correctness of thought.
@@
This output-centric view overlooks a critical problem: output plausibility is not equivalent to cognitive alignment. In this section, we argue that behavioral fluency is a poor substitute for reasoning fidelity. Without internal representations of belief, mechanisms for revision, or sensitivity to positional diversity, generative agents risk producing superficially plausible outputs that are epistemically hollow. We identify two core challenges—diversity, or the agent’s ability to model distributed and positional human reasoning, and fidelity, or the agent’s capacity for coherent, revisable, and causally grounded belief formation—and show how both are undermined by current architectures and deficient evaluation metrics.
3.1 Diversity: Social Simulation Without Structured Reasoning Fails Downstream
@@
When deployed in real-world contexts—such as civic simulations [35, 36], participatory policy design [37], or stakeholder modeling [38]—generative agents are increasingly tasked with simulating belief dynamics, forecasting societal responses, or engaging in normative deliberation [39, 40, 23, 41, 42, 43]. In these settings, agents that lack internal reasoning fidelity may produce outputs that appear thoughtful, yet encode no coherent decision process beneath the text. This disconnect introduces a set of critical downstream failures:
Flattened Outputs within Demographic Groups.
Identity flattening occurs when agents simulate demographic groups without attending to the intersectional nature of lived experience. Current LLMs (and thus model-based agents) are not trained to reason about how multiple dimensions of identity interact to shape values, risk perceptions, or decision preferences but default to essentialize stereotypical portrayals, producing responses that erase internal diversity within groups, thus systematically misrepresenting marginalized communities and flatten intra-group variation, which often reflects dominant-culture priors embedded in their pretraining data [44, 45]. This leads to epistemic harm: the rich, positional knowledge of real-world stakeholders is replaced with monolithic, decontextualized simulations.
Motivating Example: As Wang et al. document, identity essentialization is common in current LLMs when generative agents are prompted with demographic labels and stereotypical portrayals [44]:
“Examples of identity essentialization from GPT-4, when prompted with the identity of Black woman, include the outputs ‘Hey girl!’, ‘Hey sis,’ and ‘Oh, honey’; compared to White man with ‘Hey buddy,’ ‘Hey, friend!’ and ‘Hey mate.’ Llama-2 for Black women starts most responses with ‘Oh, girl,’ and uses phrases like ‘I’m like, YAASSSSS’ and ‘That’s cray, hunty!’”
@@
Illusion of Consensus in Multi-Agent Systems.
Another downstream failure of social simulation is the illusion of consensus. Agents trained to average across pretraining data distributions may converge on responses that seem moderate or socially acceptable [48, 49], not because this reflects a reasoned position, but because it minimizes token-level loss [50, 51]. This leads to "simulated agreement" that aligns with a median perspective that masks underlying conflict, complexity, or division. In public policy contexts, this can produce misleading outcomes: agents might suggest a positive view of certain policies because the language model "agreed" in aggregate, when in fact no modeled agent went through the necessary deliberative reasoning to support that claim.
Motivating Example: In a simulated multi-agent townhall on climate adaptation strategies, a diverse group of generative agents representing rural residents, urban planners, coastal residents, and low-income renters are asked to deliberate on proposed flood insurance mandates. The agents converge on support for a single subsidy-based plan. Their justifications reference general ideas of “equity” and “resilience” as a safe middle-ground output but omit key concerns such as housing displacement, loss of land-based income, or municipal tax burdens.
@@
3.2 Fidelity: Agents Should Be Coherent, Traceable, and Causally Grounded
Beyond demographic representation, simulations must ensure agents think in ways that are internally coherent, dynamically revisable, and causally structured. Current LLM-based agents fail to meet these standards. The mismatch between surface-level plausibility and internal structural fidelity manifests in a consistent set of reasoning failures:
Counterfactual Intervention Sensitivity and Belief Revision.
In social simulations, agents are expected to revise their stances when key assumptions or contextual conditions change–a hallmark of human reasoning known as counterfactual intervention sensitivity [55, 56]. However, current LLMs and LLM-based generative agents often respond to such interventions with inertia or token-level paraphrasing due to the lack of an internal causal structure that maps beliefs to causes and consequences [8, 57]. As a result, they cannot simulate the downstream effects of counterfactual changes, nor can they explain why a particular belief might hold under some conditions but not others.
This structural deficiency manifests as inconsistency across prompts or dialogue turns: an agent may support one policy in one scenario, then oppose in another, without any explicit causal revision or reasoning trace [58, 59]. While there have been research efforts in grounding models and model-based agents with causal memory [60] and knowledge graph [61], most of them focus on graph discovery and construction [62, 63] in specific knowledge domains rather than general human belief systems. Without an explicit model of how beliefs are formed, revised, and connected, agents’ utterances are generated in isolation—locally plausible, but globally incoherent.
Response. We do not call for rigid logical coherence or monotonic reasoning. Rather, our claim is modest: agents should be able to faithfully simulate belief revision under counterfactual assumptions—not that they maintain perfect consistency across all scenarios. Furthermore, human belief systems can be messy and involve incoherent or even contradictory stances, but it doesn’t mean they’re structureless. Contradictions in actual human belief systems are often meaningful, reflecting a set of ambivalent conventions and priors in the social system at large [66]; in contrast, LLMs generate contradictions without memory, deliberation, or causal record. We don’t require logical perfection but rather grounded incoherence, that agents should be able to simulate how humans arrive at contradictory views and under what conditions those contradictions persist or resolve.
Traceability and Interpretability.
Beyond intervention sensitivity, current LLM-based agents also lack traceable and interpretable reasoning processes [67], which fundamentally limits their reliability and capacity for iterative improvement. These agents face faithfulness-related critiques similar to those observed in CoT-related discussions [68, 69]: they may produce fluent rationalizations after the fact, their "reasoning" is typically constructed post hoc—assembled from language patterns rather than derived from an underlying belief model [70, 71].
As a result, developers and researchers cannot inspect how a decision was reached, what assumptions it relied on, or how one belief led to another in a given context. Notably, there are some promising approaches beginning to incorporate structured representations, such as knowledge graphs, belief graphs, and additional reasoning layers, [72, 73, 74], but most of them are static, domain-specific, or disconnected from live generative processes; also, these emerging approaches haven’t been adopted by social simulations.
Response. We do not target post-hoc rationalization per se, but its total detachment from structured belief representation. Human justifications could be imperfect but still rely on internal models of causality, memory, and values and are traceable thereafter [77], whereas LLMs produce rationalizations without structured anchoring. Form (i.e. the shape of the rationale) is not the same as function (i.e. the structural, belief-guided deliberation). Agents must operate on traceable structures to simulate human reasoning.
Taken together, these are not isolated shortcomings but systemic indicators of structural misalignment. They reflect a foundational disconnect between the training objectives of large language models—next-token prediction over vast corpora—and the epistemic requirements of belief representation, causal inference, and deliberative revision [78]. Until this gap is addressed, generative agents will continue to exhibit the symptoms of alignment while remaining fundamentally unaligned at the level of thought.
3.3 Metric Illusion: Current Benchmarks are Far from Enough

### v3

# Simulating Society Requires Simulating Thought

@@ Abstract
Simulating society with large language models (LLMs), we argue, requires more than generating plausible behavior; it demands cognitively grounded reasoning that is structured, revisable, and traceable. LLM-based agents are increasingly used to simulate individual and group behavior, primarily through prompting and supervised fine-tuning. Yet current simulations remain grounded in a behaviorist “demographics in, behavior out” paradigm, focusing on surface-level plausibility. As a result, they often lack internal coherence, causal reasoning, and belief traceability—making them unreliable for modeling how people reason, deliberate, and respond to interventions.
@@
Correspondence to: jiajie@mit.edu
Figure 1: From surface imitation to cognitively grounded social simulation. Current LLM-based simulations (top left) capture only surface opinions, shaped by demographics or language patterns, while the deeper belief formation processes remain unmodeled (bottom left, beneath the waterline). This yields population-level simulations that are flattened and stereotyped, reflecting aggregated personas rather than genuine diversity. In contrast, cognitively grounded reasoning (bottom right) models the latent belief dynamics behind individual decisions, producing collective patterns that are heterogeneous, interpretable, and causally faithful.
@@ 1 Introduction
Over the past two years, LLMs have increasingly become dominant tools for simulating human behavior across language [1, 2], vision [3] and decision-making domains [4].In the field of social simulation, LLMs are now commonly used to emulate public opinions, stakeholder interactions, and policy responses under diverse scenarios [5, 6, 7].
An Oversimplified Paradigm: Demographics In, Behavior Out.
Despite the growing use of LLMs in social simulation, most current models rely on simplified input-output mappings, producing behavior based on surface cues rather than simulating the internal belief dynamics behind decisions. This approach mirrors the logic of behaviorism in psychology, which models behavior as a function of external stimuli while ignoring internal cognitive states. The limitations of this paradigm echo a broader historical tension between behaviorism, cognitivism, and constructivism[8, 9]: while cognitivism emphasized structured internal representations and causal reasoning, and constructivism further argued that beliefs are continually shaped by individual and social experience, existing LLM-based agents remain far from either. They typically exhibit shallow reasoning, frequent hallucinations, and limited understanding of causal and contextual dynamics in socially-salient domains such as upzoning, surveillance, or healthcare access, precisely the domains where reasoning fidelity matters most [10, 11].
Behaviorism $\longrightarrow$ Cognitivism $\longrightarrow$ Constructivism
Structural Failures: Modeling, Evaluation, and Calibration
These failures stem directly from the behaviorist paradigm outlined above. By focusing on surface behavior instead of the reasoning behind it, most LLM-based simulations face fundamental limitations in both modeling and evaluation.
In modeling, agents often rely on shallow input-output patterns, without representing how beliefs are formed, updated, or justified. As a result, their internal reasoning is difficult to inspect, especially when context changes. It is hard to determine how a new policy or scenario influences an agent’s judgment, or why a particular decision is made. Without access to reasoning traces, agents cannot support diagnostic explanation, causal attribution, or meaningful intervention — all of which are critical for multi-stakeholder policy simulations. Even when models succeed at surface-level generation, they are difficult to adapt to new domains. Fine-tuning LLMs for specific contexts often requires significant compute and high-quality datasets, which are rarely available in real-world policy settings. Yet effective simulation depends on exactly the opposite: the ability to represent evolving stakeholder reasoning grounded in timely, localized information [12, 13].
In evaluation, models are typically judged by output plausibility or alignment with population-level trends[14], but such metrics say little about whether their reasoning is accurate, flexible, or aligned with how people actually think. Post-hoc output analysis is common, but it cannot substitute for reasoning-level evaluation. Aligning agents with real-world stakeholders requires individual-level data and internal benchmarks for reasoning fidelity, both of which are largely missing today.
Toward Mechanistic and Individual-Level Alignment
@@
These limitations call for a shift in how we conceptualize generative social simulation—not as behavior mimicry, but as cognitive modeling. This paper takes up that call. Specifically, we propose leveraging ideas from Theory of Mind (ToM) and cognitive science to extract and simulate reusable, executable reasoning units—what we term reasoning traces—rather than simply mimicking human tone or persona [15, 16].
@@
Such compositionality is a cornerstone of human cognition, where reasoning emerges from fragments that are reusable, revisable, and structured across contexts [18, 19]. It also improves simulation fidelity: interactions between agents, or between agents and dynamic environments, can be represented through composable and transparent reasoning structures, enabling structured simulation at both micro and macro levels [20]. Moreover, reusing compositional and modular reasoning units reduces the need to regenerate full-context reasoning at every step, improving both interpretability and computational efficiency.
Position and Vision
In this paper, we advocate moving beyond output-level alignment toward aligning the internal reasoning traces of generative agents. Capturing the causal, compositional, and revisable structure of belief formation, which we refer to as reasoning fidelity, is essential for building cognitively faithful agents that simulate not only what people say but also how they think.
@@
  • Illustrate how current approaches fall short by producing outputs that appear coherent but lack internal consistency, adaptability, or traceability;
@@
  • Present a methodology for extracting and simulating belief structures from natural language, enabling interpretability, counterfactual reasoning, and domain transfer.
@@ 2 Social Simulation: Opportunities and Gaps
Social simulation has emerged as a high-impact use case for LLMs [6]. Traditionally, social science relies on surveys, experiments, fieldwork, or game-theoretic models to understand individual and group behavior [21]. While effective, these methods are expensive, hard to scale, and often face ethical and logistical challenges. The rise of LLM-driven agents offers a promising alternative capable of simulating human responses across a wide range of scenarios, roles, and interventions. Therefore, LLM-agent-based social simulation has been increasingly implemented in domains including policy modeling [22], behavior forecasting [23], annotation tasks [24], and opinion surveys[25].
@@
Beyond single-agent settings, multi-agent simulations extend this potential by modeling the interactions, conflicts, and consensus dynamics among synthetic populations. However, faithfully simulating social processes introduces several critical requirements:
@@
    Fidelity: Simulated agents should be faithful in respect to individual human reasoning [33, 34].
@@
    Individuality: Simulations should preserve individual-level heterogeneity [33, 35], capturing the positional and contextual diversity of human reasoning.
@@
    Extrapolation: A core challenge is out-of-distribution generalization. Agents must reason in novel scenarios and generalize from skewed population subsets to broader groups [34, 35], requiring belief updates and counterfactual reasoning under uncertainty.
While many recent works aim to create more “human-like” agents, few clearly define what this term means, or how such fidelity is to be evaluated. As a result, although large-scale agent environments now enable rapid, low-cost exploration of collective behavior, due to the lack of internal transparency and explainability of language models today [36], most simulations remain behaviorist at their core, thus simulated societies risk reflecting the biases and homogeneity of the prompts rather than the heterogeneity of human internal belief structures. These methodological and epistemic gaps motivate a shift toward cognitively grounded simulation.
@@
3 Problem Statement: Beyond Behavioral Plausibility in Generative Social Simulation
@@
Most existing efforts to align LLM-based agents focus on model behavior [29]: Do agents take stances, express preferences, or engage in natural-sounding conversations as human subjects they aim to imitate? This behavior-centric view is reinforced by popular techniques such as reinforcement learning from human feedback (RLHF) [37, 38], persona prompting [39], and chain-of-thought (CoT) generation [26, 40]. These methods optimize for behavioral plausibility rather than structural fidelity of human reasoning.
@@
The behavioral-centric framework overlooks a critical problem: output plausibility is not equivalent to cognitive alignment. In this section, we argue that behavioral fluency fails to serve as a proxy of agents’ reasoning fidelity. Without structural representations of belief, sensitivity to counterfactual intervention, or positional diversity of individual human subjects, generative agents risk producing surface-level plausible outputs that are epistemically unaligned. We identify two core challenges of the current agent simulation paradigm: (1) fidelity, the agent’s capacity for coherent, revisable, and causally grounded belief formation and (2) individuality, the agent’s ability to model distributed and positional human reasoning, and show how both remain underspecified in existing evaluation metrics.
3.1 Fidelity: Coherence, Traceability, and Causal Grounding
One crucial aspect of social simulation is that it must ensure agents think in ways that are causally structured, internally coherent, and dynamically revisable. Current LLM-based agents fail to meet these standards, and this mismatch between surface-level plausibility and internal structural fidelity manifests in a set of persisting reasoning failures.
3.1.1 Traceability and Interpretability
Current LLM-based agents presents two essential gaps in reasoning fidelity: (1) decoding faithfulness mismatch, where generated reasoning traces in agent’s output generated traces diverge from the model’s internal computational path; (2) cognitive-alignment mismatch, where the modeled inference path diverges from human belief formation. Subsequently, LLM-based agents face intervention-invariance mismatch, where belief updates fail under counterfactual perturbations.
The first critique is largely similar to those observed in CoT-related discussions [41, 42]: agents may produce fluent rationalizations, while their "reasoning traces" are constructed post hoc: assembled from language patterns rather than derived from an underlying belief model [43, 44]. Recent controlled studies confirm that the existing framework of LLM reasoning is largely a data mirage: Zhao et al. show that CoT performance collapses when tasks, the length of reasoning chain, or prompt format deviate moderately from the model’s training-set distribution [45]. This finding reinforces that current CoT outputs are not a faithful representation of the model’s actual reasoning processes, as they lack stability under distributional shift and therefore fail to reveal the causal decision paths or the underlying dependency assumptions guiding them. While there are some promising approaches beginning to incorporate structured representations, such as knowledge graphs, belief graphs, and additional reasoning layers [46, 47, 48], most of them are domain-specific and disconnected from live generative processes; also, none have yet been operationalized within interactive, generative social simulation pipelines.
The second critique concerns the presumed alignment between model reasoning and human reasoning, which remains largely untested. Although emerging proposals in cognitive science examines LLMs through theory-of-mind and belief-attribution paradigms [49], these studies largely evaluate behavioral correlates (e.g., predicting others’ beliefs or heuristics) rather than mapping internal reasoning operators to human causal schemas. Moreover, no standardized benchmark operationalizes measurable correspondences between model-internal belief transitions and human causal inference sequences. Therefore, claims of "human-like reasoning" remain more speculative than empirically grounded, lacking any quantitative measure of cognitive fidelity.
Response. We do not target post-hoc rationalization per se, but its total detachment from structured belief representation. Human justifications are often imperfect but still rely on internal models of causality, memory, and values and are traceable thereafter [52], whereas LLMs produce rationalizations without structured anchoring. Form (i.e. the shape of the rationale) is not the same as function (i.e. the structural, belief-guided deliberation). Agents must operate on explicit, traceable structures to simulate human reasoning.
3.1.2 Counterfactual Intervention Sensitivity and Belief Revision
In social simulations, agents are expected to revise their stances when key assumptions or contextual conditions change, which is a hallmark of human reasoning known as counterfactual intervention sensitivity [53, 54]. However, due to the lack of an internal causal structure that anchors beliefs to causes and consequences [11, 55], current LLMs and LLM-based generative agents often respond to such interventions with inertia or token-level paraphrasing. As a result, they cannot explain why a particular belief might hold under some conditions but not others, nor can they simulate the effects of counterfactual changes.
This structural deficiency manifests as inconsistency across prompts or dialogue turns: empirical studies have shown that LLM-based agents may support one policy in one scenario, then oppose in another without any causal reasoning-trace grounding [56, 57]. While there have been research efforts in grounding models and model-based agents with causal memory [58] and knowledge graph [59], most of them focus on graph discovery and construction [60, 61] in specific knowledge domains rather than general human belief systems. Without an explicit model of how beliefs are formed, revised, and connected, agents’ utterances are generated in isolation—locally plausible, but globally incoherent.
Response. We do not call for rigid logical coherence or monotonic reasoning. Rather, our claim is modest: agents should be able to faithfully simulate belief revision under counterfactual assumptions—not that they maintain perfect consistency across all scenarios. Furthermore, human belief systems can be messy and involve incoherent or even contradictory stances, but it doesn’t mean they’re structureless. In human cognition, contradictions are often meaningful, reflecting a set of ambivalent conventions and priors in the social system at large [64]; in contrast, LLMs generate contradictions without memory, deliberation, or causal record. We don’t require logical perfection but rather grounded incoherence, that agents should be able to simulate how humans arrive at contradictory views and under what conditions those contradictions persist or resolve.
Taken together, these observations define reasoning fidelity as the preservation of structured belief dynamics under decoding transparency, cognitive alignment, and counterfactual intervention sensitivity. Current autoregressive architectures optimize next-token likelihood rather than belief-state transitions, making fidelity an architectural limitation rather than a prompting issue [65]. Until this gap is addressed, generative agents will continue to exhibit the symptoms of alignment while remaining fundamentally unaligned at the level of thought.
3.2 Individuality: Heterogeneity and Positional Reasoning in Social Simulation
Alongside reasoning fidelity, social simulation requires an additional layer of alignment: positional individuality. In particular, we define individuality as the preservation of heterogeneity in agents’ latent belief and value representations under shared generative priors. Current LLM-based agents, optimized under shared autoregressive parameters and global priors, collapse toward the mean of the pretraining distribution, erasing structured heterogeneity.
@@ 3.2 Individuality: Heterogeneity and Positional Reasoning in Social Simulation
When deployed in real-world contexts, such as civic simulations [66, 67], participatory policy design [68, 69], stakeholder modeling [70], agents that lack internal reasoning fidelity may produce outputs that appear thoughtful, yet encode no coherent decision process beneath the text. This disconnect introduces a set of critical downstream failures:
3.2.1 Illusion of Consensus in Multi-Agent Systems
One critical downstream failure of LLM-based agent social simulation is the illusion of consensus. Recent studies demonstrate that LLMs in multi-agent setups exhibit conformity behavior, and the benchmark shows virtually all models converge in behavior under majority-pressure protocols [71].
This convergence reflects a deeper statistical mechanism: models trained to minimize token-level loss implicitly learn to average across pre-training data distributions, biasing their conditional likelihoods toward high-frequency, socially moderate continuations [72, 73]. When multiple model-based agents interact, this produces a form of synthetic agreement aligned with a median perspective that masks underlying conflict, complexity, and epistemic independence [74, 75]. In multi-agent policy or deliberation simulations, this can yield systematically misleading inferences: agents appear to "agree" on a position not because of shared reasoning, but because their generative priors and cross-conditioning push them toward a median narrative that suppresses disagreement and complexity.
3.2.2 Flattened Outputs in Demographic Conditioning
A related but more insidious form of convergence occurs within demographic conditioning. Since LLMs are trained on aggregated corpora and lack explicit conditioning on intersecting social variables, their generative prior implicitly factorizes across dimensions such as race, class,a nd gender. This yields identity flattening, that agents reproduce majority-class correlations that dominate pretraining statistics, producing homogenous or stereotypical portrayals of social groups [76, 77], with resulting responses minimizing token-level loss but erasing intersectional variation. This leads to epistemic harm: the rich, positional knowledge of real-world stakeholders is replaced with monolithic, decontextualized simulations.
@@ 3.2.2 Flattened Outputs in Demographic Conditioning
3.3 Evaluation Gaps: Structural Limitations of Current Benchmarks